0.2g of anhydrous organic acid gave on combustion 0.04g water and 0.195gof carbon di oxide. The acid is found to be dibasic and 0.5g of its silver salt leaves on ignition 0.355g silver. The molecular formula of the compound is

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Pallavi,

0.5 g silver salt-------------0.355 g silver--------------0.145g

R(COOAg)2                            Ag                          R(COO-)2

0.355/108----no of moles of Ag

0.355/216----no of moles of   R(COO-)2

Mol. Wt. of   R(COO-)2 is 0.145/(0.355/216) = 88

Mol. Wt. of  R(COOH)2 is 88+2 = 90

0.2 g of Acid------0.2/90= 0.002222 moles

0.002222 moles acid-------- 0.04*2/18 = 0.004444 moles H--------0.195/44 = 0.00444 moles of C

1 mole acid------2 moles H-------2 moles C

(COOH)2 is the Molecular formula of the Organic acid.